5. a. Write out the first 4 partial sums of the series -(-1). b. Is In-,(-1) a telescoping series? Explain why or why not.
Answer #1

Sla) 9 (-” no sum first four partial Si = (-10° = 60°+(-0 = 1-1=0 Sz = 61,9+ (-1) + (-1)! = 1-1+1 = 1 sy - G10+(-1)+(-0,2+(